OpenStudy (anonymous):

U^-3=1/125 U=? @GretaKnows

myininaya (myininaya):

hint: 5^3=125

myininaya (myininaya):

one more hint: \[\frac{1}{a^n}=a^{-n}\]

OpenStudy (anonymous):

Im sorry but im lost??=/

myininaya (myininaya):

here is an example: \[u^{-3}=\frac{1}{8} \\ u^{-3}=\frac{1}{2^3} \\ u^{-3}=2^{-3} \\ u=2\]

OpenStudy (anonymous):

Ok thanks I got 5?

myininaya (myininaya):

sounds great